### v1

@@ 1 Introduction
Significant efforts have been made to accelerate attention computation through software fusion techniques on datacenter-grade hardware. For cloud servers, multi-core parallelism Shoeybi et al. (2019); Rasley et al. (2020); Narayanan et al. (2021); Kwon et al. (2023); Liu et al. (2023); Cho et al. (2024) and efficient utilization of on-chip SRAM in GPUs Kirk et al. (2007) are employed to enhance performance. FlashAttention Dao et al. (2022); Dao (2023); Shah et al. (2024); dao ; Hong et al. (2023) related methods design I/O-aware exact attention speedup algorithms, leveraging GPU CUDA cores and on-chip SRAM to minimize access to High Bandwidth Memory (HBM), saving memory and reducing runtime. FuseMax Nayak et al. (2024) uses Einsums and a spatial array accelerator, employing ping-pong scheduling to overlap MatMul and softmax operations. While FlashAttention-3 and FuseMax parallelize MatMul and softmax on multi-core architectures, these cloud-based acceleration methods cannot be applied on resource-constrained edge accelerators directly, which have a limited number of processing units and on-chip memory.
@@ 2 Related Work
Exact Attention Acceleration Methods: In cloud environments, exact acceleration methods Dao et al. (2022); Dao (2023); Shah et al. (2024); dao ; Hong et al. (2023); Patel et al. (2023); Nayak et al. (2024) leverage parallel computation on multi-core architectures to speed up attention mechanism. For instance, FlashAttention and FlashAttention-2 optimize dataflow for attention computation on NVIDIA A100 GPUs by dividing query, key and value inputs into smaller tiles and loading them from high-bandwidth memory (HBM) to on-chip SRAM, reducing data movement for large intermediate outputs and exploiting GPU CUDA core parallelism. FlashAttention-3 further enhances parallelism using ping-pong scheduling to overlap MatMul and softmax operations within warp groups on NVIDIA H100 GPUs. In contrast, FuseMax uses Einsums to design a spatial array accelerator for fused attention calculation in cloud environments, overlapping MatMul and softmax computations and improving spatial PE array utilization.
@@ 4.2 MAS-Attention Tiling Scheme
The Genetic Algorithm evolves a population of tiling configurations through selection, crossover, and mutation, iteratively refining the configurations to optimize execution cycles and power consumption. MCTS, on the other hand, builds a search tree where each node represents a potential tiling configuration, balancing exploration of new configurations with exploitation of high-performing ones to converge toward optimal or near-optimal solutions. On the DaVinci DNN Accelerator, Grid Search systematically evaluates all possible configurations, leveraging its compatibility with the hardware’s structured memory model. These algorithms iteratively assess various tiling configurations, simulating different tile shapes and sizes to determine those that optimize execution cycles and minimize power consumption.
@@ 5.1 Experimental Setup
Simulation and Modeling Tools: To simulate our experiments, we employed Timeloop Parashar et al. (2019) and Accelergy Wu et al. (2019) to measure the latency and energy consumption, also we modified TileFlow Zheng et al. (2023) to define the edge spatial accelerator, software mapping for attention inference, and search space exploration. The custom edge hardware architecture designed for simulation operates at a frequency of 3.75GHz and features 16nm technology, two cores each containing a MAC and a VEC unit, and a hierarchical memory system as depicted in Figure 4. The designed DRAM has a bandwidth of 30GB/sec and a total size of 6GB. The L1 cache has connection to DRAM and L0 register file and has a storage of 5MB . The Processing Elements (PEs) in MAC and VEC units, organized in 16x16 and 256 mesh respectively, have access to L0 register file. These parameters for the hardware architecture were determined after various stress tests of the hardware. Our simulations were conducted on a system equipped with an Intel(R) Xeon(R) Gold 6140 CPU @ 2.30GHz, utilizing a single thread in execution. Additionally, we evaluated our algorithm on a real hardware, specifically Huawei MatePad Pro 13.2 with DaVinci DNN Accelerator, to validate its practical applicability and performance.
@@
Workloads: The workload for our experiments focuses on the inference of attention layers in various transformer-based networks, including different variants of BERT Devlin et al. (2018), Llama Touvron et al. (2023), T5 Raffel et al. (2020), ViT Dosovitskiy et al. (2020), and XLM Lample & Conneau (2019), as detailed in Table 1. We selected a diverse set of networks with varying attention layer dimensions to ensure a comprehensive evaluation. Additionally, each workload in our experiments undergoes a rigorous golden data check for all methods, including our proposed approach, ensuring that all methods pass this validation.
@@ 5.2.1 Analysis on Simulated Hardware
Table 2 presents a detailed analysis of execution cycles and speedup ratios for MAS-Attention compared to other methods across all tested networks. The data highlights that MAS-Attention consistently achieves superior performance, with speedup factors ranging from 1.02×1.02$\times$1.02 × to 8.50×8.50$\times$8.50 × over Layer-Wise, and up to 5.25×5.25$\times$5.25 × over Soft-Pipe, 2.75×2.75$\times$2.75 × over FLAT, and 2.00×2.$0$ over TileFlow methods. The geometric mean of these speedup values—5.09×5.09$\times$5.09 ×, 2.78×2.78$\times$2.78 ×, 1.70×1.70$\times$1.70 ×, and 1.31×1.31$\times$1.31 × respectively—demonstrates MAS-Attention’s overall efficiency in reducing execution time. This substantial performance improvement underscores MAS-Attention’s effectiveness as an advanced solution for optimizing computational efficiency in attention mechanisms.
@@
Layer-Wise Soft-Pipe FLAT TileFlow MAS-Attention Layer-Wise Soft-Pipe FLAT TileFlow
BERT-Base & T5-Base 3.637 2.064 1.573 0.799 0.786 4.63 2.63 2.00 1.02
BERT-Large & T5-Large 5.505 2.753 1.835 1.311 1.049 5.25 2.63 1.75 1.25
BERT-Small 2.753 1.376 0.918 0.655 0.524 5.25 2.63 1.75 1.25
Llama3-8B & T5-3B (T5-XL) 12.845 8.389 4.719 5.243 4.194 3.06 2.00 1.13 1.25
T5-Mini & T5-Small 2.228 1.180 0.721 0.328 0.262 8.50 4.50 2.75 1.25
ViT-B/14 0.612 0.381 0.266 0.263 0.151 4.06 2.53 1.77 1.75
ViT-L/14 1.242 0.508 0.354 0.351 0.201 6.19 2.53 1.77 1.75
ViT-H/14 1.355 0.558 0.405 0.439 0.251 5.40 2.23 1.61 1.75
ViT-B/16 1.081 0.590 0.426 0.249 0.197 5.50 3.00 2.17 1.27
ViT-L/16 1.311 0.786 0.524 0.332 0.262 5.00 3.00 2.00 1.27
ViT-H/16 1.376 0.852 0.590 0.414 0.328 4.20 2.60 1.80 1.26
XLM 4.194 2.097 1.180 1.311 1.049 4.00 2.00 1.13 1.25
Geometric Mean - - - - - 5.09x 2.78x 1.70x 1.31x
@@ 5.3 Power and Energy Analysis
Table 3 presents a comprehensive analysis of energy consumption and savings achieved by MAS-Attention compared to other methods across various networks. The data reveals that MAS-Attention consistently demonstrates significant energy efficiency improvements, with savings ranging from $1.98\%$ to $75.00\%$ compared to Layer-Wise, $36.83\%$ to $66.08\%$ compared to Soft-Pipe, $0.02\%$ to $54.03\%$ compared to FLAT, and $36.83\%$ to $65.05$ compared to TileFlow. The geometric mean of these savings—$52.97\%$, $63.07\%$, $18.55\%$, and $53.16\%$ respectively—highlights MAS-Attention’s overall effectiveness in reducing energy consumption. This substantial reduction underscores MAS-Attention’s potential for optimizing energy efficiency in resource limited environments.
@@
Network Name Energy Consumption ($10^{9}$ nJ) Energy Savings (MAS-Attention vs. Others)
Layer-Wise Soft-Pipe FLAT TileFlow MAS-Attention Layer-Wise Soft-Pipe FLAT TileFlow
BERT-base & T5-Base 37.208 49.607 12.656 27.598 12.405 66.67% 75.00% 1.98% 55.05%
BERT-large & T5-Large 28.105 65.672 21.112 38.065 16.944 39.69% 74.20% 19.75% 55.49%
BERT-small 20.218 24.336 10.556 19.032 8.359 58.65% 65.64% 20.80% 56.08%
Llama3-8B & T5-3B (T5-XL) 179.309 186.463 63.252 147.502 63.241 64.73% 66.08% 0.02% 57.12%
T5-Mini & T5-Small 12.434 11.269 8.744 7.512 4.746 61.83% 57.90% 45.71% 36.83%
ViT-B/14 3.720 7.376 2.803 4.136 1.903 48.87% 74.21% 32.11% 54.00%
ViT-L/14 5.539 7.335 5.648 7.428 2.596 53.13% 64.61% 54.03% 65.05%
ViT-H/14 6.585 9.120 4.741 6.783 3.162 51.98% 65.34% 33.27% 53.38%
ViT-B/16 5.323 5.828 3.350 7.119 3.239 39.16% 44.42% 3.34% 54.49%
ViT-L/16 9.403 6.984 6.316 9.402 4.218 55.14% 39.61% 33.21% 55.14%
ViT-H/16 11.160 15.414 6.803 11.475 5.156 53.81% 66.55% 24.22% 55.09%
XLM-Base 35.786 46.485 15.813 36.876 15.584 56.45% 66.47% 1.45% 57.74%
Geometric Mean - - - - - 52.97% 63.07% 18.55% 53.16%
Note: Based on some literature studies, “nJ” (nanojoules) is used as the unit for energy consumption reported by Accelergy.

### v2

@@ 1 Introduction
Significant efforts have been made to accelerate attention computation through software fusion techniques on datacenter-grade hardware. For cloud servers, multi-core parallelism Shoeybi et al. (2019); Rasley et al. (2020); Narayanan et al. (2021); Kwon et al. (2023); Liu et al. (2023); Cho et al. (2024) and efficient utilization of on-chip SRAM in GPUs Kirk et al. (2007) are employed to enhance performance. FlashAttention Dao et al. (2022); Dao (2023); Shah et al. (2024); dao ; Hong et al. (2023) related methods design I/O-aware exact attention speedup algorithms, leveraging GPU CUDA cores and on-chip SRAM to minimize access to High Bandwidth Memory (HBM), saving memory and reducing runtime. FuseMax Nayak et al. (2024) uses Einsums and a spatial array accelerator, employing ping-pong scheduling to overlap MatMul and softmax operations. While FlashAttention-3 Shah et al. (2024) parallelizes MatMul and softmax on multi-core architectures, these cloud-based acceleration methods do not directly apply to resource-constrained edge accelerators, where there are limited number of processing units and on-chip memory.
@@ 2 Related Work
Exact Attention Acceleration Methods: In cloud environments, exact acceleration methods Dao et al. (2022); Dao (2023); Shah et al. (2024); dao ; Hong et al. (2023); Patel et al. (2023) leverage parallel computation on multi-core architectures to speed up attention mechanism. For instance, FlashAttention and FlashAttention-2 optimize dataflow for attention computation on NVIDIA A100 GPUs by dividing query, key and value inputs into smaller tiles and loading them from high-bandwidth memory (HBM) to on-chip SRAM, reducing data movement for large intermediate outputs and exploiting GPU CUDA core parallelism. FlashAttention-3 further enhances parallelism using ping-pong scheduling to overlap MatMul and softmax operations within warp groups on NVIDIA H100 GPUs. FuseMax Nayak et al. (2024) leverages Einsums to implement fused attention computation on a spatial array accelerator, overlapping MatMul and softmax operations to enhance spatial PE array utilization.
@@ 4.2 MAS-Attention Tiling Scheme
We use MCTS to optimize tiling factors. At each step, MCTS selects a loop and assigns a tiling factor based on the number of iterations the loop will execute, updating constraints and passing them to the next untiled loop. Once all tiling factors are determined, a complete fusion mapping is produced as an analysis tree where each node corresponds to a tile, which is then evaluated. The results of each evaluation are fed back to MCTS to update the upper confidence bounds (UCB), guiding subsequent searches. Genetic algorithm (GA) then aims to find optimal compute ordering in the analysis tree based on the found tiling factors, refining performance across different analysis trees. GA generates a population of analysis trees, applies crossover and mutation, and evaluates each tree using the tiling factors. Through repeated iterations, the best analysis tree is selected as the optimal fusion dataflow.
On the DaVinci DNN Accelerator, Grid Search systematically evaluates all possible configurations, leveraging its compatibility with the hardware’s structured memory model. These algorithms iteratively assess various tiling configurations, simulating different tile shapes and sizes to determine those that optimize execution cycles and minimize power consumption.
@@ 5.1 Experimental Setup
Simulation and Modeling Tools: To simulate our experiments, we employed Timeloop Parashar et al. (2019) and Accelergy Wu et al. (2019) to measure the latency and energy consumption, also we modified TileFlow Zheng et al. (2023) to define the edge spatial accelerator, software mapping for attention inference, and search space exploration. During the tiling and loop parameters search, MCTS generated tiling factors and GA refined compute orderings, with each candidate evaluated using Timeloop/Accelergy. The custom edge hardware architecture designed for simulation operates at a frequency of 3.75GHz and features 16nm technology, two cores each containing a MAC and a VEC unit, and a hierarchical memory system as depicted in Figure 4. The designed DRAM has a bandwidth of 30GB/sec and a total size of 6GB. The L1 cache has connection to DRAM and L0 register file and has a storage of 5MB . The Processing Elements (PEs) in MAC and VEC units, organized in 16x16 and 256 mesh respectively, have access to L0 register file. These parameters for the hardware architecture were determined after various stress tests of the hardware. Our simulations were conducted on a system equipped with an Intel(R) Xeon(R) Gold 6140 CPU @ 2.30GHz, utilizing a single thread in execution. Additionally, we evaluated our algorithm on a real hardware, Huawei MatePad Pro 13.2, to validate its practical applicability and performance. More specifically, this device is equipped with the Kirin 990 5G SoC featuring Da Vinci NPU architecture, which consists of three cores—each with a MAC unit, a Vector Unit, and dedicated on-chip memory. The NPU includes 2x Ascend Lite cores and 1x Ascend Tiny core.
@@
Workloads: The workload for our experiments focuses on the inference of attention layers in various transformer-based networks, including different variants of BERT Devlin et al. (2018), Llama Touvron et al. (2023), T5 Raffel et al. (2020), ViT Dosovitskiy et al. (2020), and XLM Lample & Conneau (2019), as detailed in Table 1. We also provide end-to-end results from deploying MAS-Attention in a real-world AI workload based on Stable Diffusion 1.5 UNet discussed further in 5.2.2. We selected a diverse set of networks with varying attention layer dimensions to ensure a comprehensive evaluation. Additionally, each workload in our experiments undergoes a rigorous golden data check for all methods, including our proposed approach, ensuring that all methods pass this validation.
@@
FuseMax (scaled down to edge device): The computation is decomposed into a sequence of 12 primitive operators based on extended einsum notation, which are executed using pipelining. The attention scores are computed as $C=QK^{T}$, and the Softmax function is implemented through a series of sub-operations, where MAC and VEC are processed in parallel. The weighted sum with $V$ is fused into the Softmax pipeline itself. All computations are fused and executed in a single pass.
@@ 5.2.1 Analysis on Simulated Hardware
Table 2 presents a detailed analysis of execution cycles and speedup ratios for MAS-Attention compared to other methods across all tested networks. The data highlights that MAS-Attention consistently achieves superior performance, with speedup factors up to 8.50×8.50$\times$8.50 × over Layer-Wise, 4.5×4.5$\times$4.5 × over Soft-Pipe, 2.75×2.75$\times$2.75 × over FLAT, 1.75×1.75$\times$1.75 × over TileFlow, and 1.47×1.47$\times$1.47 × over FuseMax methods. The geometric means of these speedup values—5.09×5.09$\times$5.09 ×, 2.78×2.78$\times$2.78 ×, 1.70×1.70$\times$1.70 ×, 1.31×1.31$\times$1.31 ×, and 1.27×1.27$\times$1.27 × respectively—demonstrate MAS-Attention’s overall efficiency in reducing execution time. This substantial performance improvement underscores MAS-Attention’s effectiveness as an advanced solution for optimizing computational efficiency in attention mechanisms.
@@
Layer-Wise Soft-Pipe FLAT TileFlow FuseMax MAS-Attention Layer-Wise Soft-Pipe FLAT TileFlow FuseMax
BERT-Base & T5-Base 3.637 2.064 1.573 0.799 0.992 0.786 4.63 2.63 2.00 1.02 1.26
BERT-Large & T5-Large 5.505 2.753 1.835 1.311 1.323 1.049 5.25 2.63 1.75 1.25 1.26
BERT-Small 2.753 1.376 0.918 0.655 0.661 0.524 5.25 2.63 1.75 1.25 1.26
Llama3-8B & T5-3B (T5-XL) 12.845 8.389 4.719 5.243 4.864 4.194 3.06 2.00 1.13 1.25 1.16
T5-Mini & T5-Small 2.228 1.180 0.721 0.328 0.384 0.262 8.50 4.50 2.75 1.25 1.47
ViT-B/14 0.612 0.381 0.266 0.263 0.196 0.151 4.06 2.53 1.77 1.75 1.30
ViT-L/14 1.242 0.508 0.354 0.351 0.262 0.201 6.19 2.53 1.77 1.75 1.30
ViT-H/14 1.355 0.558 0.405 0.439 0.318 0.251 5.40 2.23 1.61 1.75 1.27
ViT-B/16 1.081 0.590 0.426 0.249 0.259 0.197 5.50 3.00 2.17 1.27 1.32
ViT-L/16 1.311 0.786 0.524 0.332 0.346 0.262 5.00 3.00 2.00 1.27 1.32
ViT-H/16 1.376 0.852 0.590 0.414 0.419 0.328 4.20 2.60 1.80 1.26 1.28
XLM 4.194 2.097 1.180 1.311 1.216 1.049 4.00 2.00 1.13 1.25 1.16
Geometric Mean - - - - - - 5.09x 2.78x 1.70x 1.31x 1.27x
@@ 5.2.2 Analysis on Real Hardware
Additionally, to provide end-to-end experimental results, we evaluated MAS-Attention on a real-world generative AI workload, specifically a reduced UNet module of Stable Diffusion 1.5 running directly on the mobile device. This UNet contains 15 attention units, with the largest attention layer featuring 2 heads, a sequence length of 4096, and an embedding size of 64. Compared to the Layer-Wise method, MAS-Attention achieved a 29.4% runtime reduction for the largest attention unit and a 6% overall reduction in end-to-end model inference latency, further demonstrating the practical effectiveness of our proposed algorithm.
@@ 5.3 Power and Energy Analysis
Table 3 presents a comprehensive analysis of energy consumption and savings achieved by MAS-Attention compared to other methods across various networks. The data reveals that MAS-Attention consistently demonstrates significant energy consumption reductions over Layer-Wise, Soft-Pipe, FLAT, and TileFlow, with savings ranging from $39.16\%$ to $66.67\%$, $39.61\%$ to $75.00\%$, $0.02\%$ to $54.03\%$, and $36.83\%$ to $65.05\%$, respectively. The geometric mean of these savings—$52.97\%$, $63.07\%$, $18.55\%$, and $53.16\%$—highlights MAS-Attention’s overall effectiveness in reducing energy consumption. When compared to FuseMax, MAS-Attention achieves lower energy consumption for ViT-B/14, ViT-L/14, ViT-H/14, ViT-L/16, and ViT-H/16 but exhibits higher energy usage in other cases. The reason is that our objective in the search framework was to minimize latency rather than energy, although MAS-Attention can be revised to optimize other objectives. Nevertheless, MAS-Attention remains competitive in these results by maintaining a strong balance between energy efficiency and overall computational cycles.
@@
Network Name Energy Consumption ($10^{9}$ pJ) Energy Savings (MAS-Attention vs. Others)
Layer-Wise Soft-Pipe FLAT TileFlow FuseMax MAS-Attention Layer-Wise Soft-Pipe FLAT TileFlow FuseMax
BERT-base & T5-Base 37.208 49.607 12.656 27.598 10.217 12.405 66.67% 75.00% 1.98% 55.05% -21.42%
BERT-large & T5-Large 28.105 65.672 21.112 38.065 13.623 16.944 39.69% 74.20% 19.75% 55.49% -24.38%
BERT-small 20.218 24.336 10.556 19.032 6.811 8.359 58.65% 65.64% 20.80% 56.08% -22.73%
Llama3-8B & T5-3B (T5-XL) 179.309 186.463 63.252 147.502 53.401 63.241 64.73% 66.08% 0.02% 57.12% -18.43%
T5-Mini & T5-Small 12.434 11.269 8.744 7.512 3.542 4.746 61.83% 57.90% 45.71% 36.83% -33.99%
ViT-B/14 3.720 7.376 2.803 4.136 2.104 1.903 48.87% 74.21% 32.11% 54.00% 9.56%
ViT-L/14 5.539 7.335 5.648 7.428 2.805 2.596 53.13% 64.61% 54.03% 65.05% 7.45%
ViT-H/14 6.585 9.120 4.741 6.783 3.487 3.162 51.98% 65.34% 33.27% 53.38% 9.31%
ViT-B/16 5.323 5.828 3.350 7.119 3.187 3.239 39.16% 44.42% 3.34% 54.49% -1.63%
ViT-L/16 9.403 6.984 6.316 9.402 4.249 4.218 55.14% 39.61% 33.21% 55.14% 0.73%
ViT-H/16 11.160 15.414 6.803 11.475 5.278 5.156 53.81% 66.55% 24.22% 55.09% 2.31%
XLM-Base 35.786 46.485 15.813 36.876 13.350 15.584 56.45% 66.47% 1.45% 57.74% -16.77%
Geometric Mean - - - - - - 52.97% 63.07% 18.55% 53.16% -11.94%
Note: Based on some literature studies, “pJ” (picojoule) is used as the unit for energy consumption reported by Accelergy. Negative savings indicate higher energy consumption compared to the baseline.
@@ 5.4.2 DRAM Read Operations
5.5 Impact of Search Algorithms on Tiling Optimization
Figure 7 illustrates the impact of employing MCTS and GA search algorithms in optimizing tile configurations for attention workloads. For clarity, the plot proportionally reduces the number of plotted lines to approximately 2K. It becomes evident that after around 10K iterations, each algorithm consistently converges toward optimal tiling parameters. Detailed final cycle counts and corresponding energy consumption metrics upon completion of the search are comprehensively listed in Tables 2 and 3. FuseMax results in these tables and its original work were obtained via its manually selected tiling sizes for tensors on different memory levels, thus excluded from Figure 7 on search convergence.
To further underscore the efficacy of the proposed search scheme with MAS-Attention, notable cycle improvements include a 64.5× reduction for BERT-Base and T5-Base (from 50.33M to 0.78M), a 16.1× reduction for BERT-Large and T5-Large (from 16.77M to 1.04M), and a similar 16.1× improvement for BERT-Small (from 8.38M to 0.52M) as well as T5-Mini and T5-Small (from 4.19M to 0.26M). Furthermore, Vision transformer workloads demonstrate significant benefits with up to 66.2× speedup—ViT-B,L,H/14 see 49.7×/24.5×/24.6× (from 7.45M/4.91M/6.14M to 0.15M/0.20M/0.25M), ViT-B,L,H/16 show 66.2×/32.2×/32.8× (from 12.58M/8.38M/10.48M to 0.19M/0.26M/0.32M). Lastly, XLM sees a 32.2× drop (from 33.55M to 1.04M), further validating the broad applicability and robustness of the search-based optimization approach.
Figure 7: Execution cycles vs. search time (both log scale) for different attention acceleration methods, demonstrating the impact of Genetic Algorithm (GA) and Monte Carlo Tree Search (MCTS) on each algorithm’s efficiency
@@ 5.6 Limitations
Furthermore, MAS-Attention’s stream processing efficiency relies on the availability of separate compute engines for MatMul and Softmax operations, leveraging dedicated MAC and VEC units for parallel execution. Therefore, MAS-Attention remains particularly effective on architectures with distinct heterogeneous compute resources—a design choice becoming increasingly common in modern edge accelerators to optimize for energy efficiency.